# Removable singularities for Lipschitz fractional caloric functions in time varying domains

### v3

@@ 4.1. Existence of removable sets with positive $\pazocal{H}_{p_{s}}^{n+1}$ measure
[…]
@@
If #$\#(E_{k})$ is the number of cubes of $E_{k}$, the above property followed, in essence, from
(4.5) #$\#(E_{k})\,\ell_{k}^{n}=2^{k(n+1)}\,\ell_{k}^{n}=1,\;\;\forall k\geq 1.$
@@
Hence, it is clear that in order to obtain a potentially non-removable Cantor set $E$, one has to enlarge it. One way to do it (motivated by [MPrT, §6]) is as follows: let us fix $s\in(1/2,1]$ and choose what we call the non-self-intersection parameter $\delta\in\mathbb{Z}_{+}$, the minimum integer $\delta=\delta(s)\geq 2$ satisfying
$s>\frac{\log_{\delta}(\delta+1)}{2},\qquad\text{that is}\qquad\delta+1<\delta^{2s}.$
[…]
$l_{\delta}:=\frac{1-\delta\lambda_{1}}{\delta-1},\qquad J_{j}:=\big{[}(j-1)(\lambda_{1}+l_{\delta}),j\lambda_{1}+(j-1)l_{\delta}\big{]},\quad j=1,\ldots,\delta,$
we keep the following union of $\delta$ closed disjoint intervals of length $\lambda_{1}$
$T_{\delta}:=\bigcup_{j=1}^{\delta}J_{j}.$
Finally, for the remaining temporal interval $[0,1]$, we do the same splitting but in $\delta+1$ intervals of length $\lambda_{1}^{2s}$. That is, we name
$\widetilde{l}_{\delta}:=\frac{1-(\delta+1)\lambda_{1}^{2s}}{\delta},\qquad\widetilde{J}_{j}:=\big{[}(j-1)(\lambda_{1}^{2s}+\widetilde{l}_{\delta}),j\lambda_{1}^{2s}+(j-1)\widetilde{l}_{\delta}\big{]},\quad j=1,\ldots,\delta+1.$
Now we keep the union of $\delta+1$ closed disjoint intervals of length $\lambda_{1}^{2s}$
$\widetilde{T}_{\delta}:=\bigcup_{j=1}^{\delta+1}\widetilde{J}_{j}.$
From the above, we define the first generation of the Cantor set as $E_{1,p_{s}}:=(T_{\delta})^{n}\times\widetilde{T}_{\delta}$, that is conformed by $(\delta+1)\delta^{n}$ disjoint $s$-parabolic cubes (see Figure 1). This procedure continues inductively, i.e. the next generation $E_{2,p_{s}}$ will be the family of $(\delta+1)^{2}\delta^{2n}$ disjoint $s$-parabolic cubes of side length $\ell_{2}:=\lambda_{1}\lambda_{2}$, $\lambda_{2}<1/\delta$, obtained from applying the previous construction to each of the cubes of $E_{1,p_{s}}$. More generally, the $k$-th generation $E_{k,p_{s}}$ will be formed by $(\delta+1)^{k}\delta^{nk}$ disjoint $s$-parabolic cubes with side length $\ell_{k}:=\lambda_{1}\cdots\lambda_{k}$, $\lambda_{k}<1/\delta$, and with locations determined by the above iterative process. We name such cubes $Q^{k}_{j}$, with $j=1,\ldots,(\delta+1)^{k}\delta^{nk}$. The resulting $s$-parabolic Cantor set is
@@
(a) Depiction of $E_{1,p_{1}}$ with $\delta=2$
(b) Depiction of $E_{1,p_{2/3}}$ with $\delta=3$.
@@
This way, the critical value becomes $((\delta+1)\delta^{n})^{-1/(n+1)}<1/\delta$. For instance, if
(4.7) $\lambda_{j}:=((\delta+1)\delta^{n})^{-1/(n+1)},\qquad\text{for every }\,j,$
@@
#$\#(E_{k,p_{s}})\cdot\ell_{k}^{n+1}=(\delta+1)^{k}\delta^{nk}\cdot\ell_{k}^{n+1}=1,\;\;\forall k\geq 1.$
[…]
(4.8) $\mu(Q)\lesssim\frac{1}{((\delta+1)\delta^{n})^{(k+1)}}=\ell_{k+1}^{n+1}\leq\ell(Q)^{n+1}$
@@
$\pazocal{H}^{n+1}_{p_{s},\varepsilon}(E)\leq\sum_{j=1}^{((\delta+1)\delta^{n})^{k}}\text{diam}_{p_{s}}(Q^{k}_{j})^{n+1}\simeq\ell_{k}^{n+1}\,((\delta+1)\delta^{n})^{k}=1.$
@@
Let us observe that as $s\to 1/2$, the value of $\delta$ grows arbitrarily. That is, as $s$ approaches the value $1/2$, the Cantor set $E_{p_{s}}$ becomes progressively more and more dense in the unit cube. This suggests that for $s=1/2$, that corresponds to a space-time usual Lipschitz condition in $\mathbb{R}^{n+1}$, the capacity one would obtain should be comparable to the Lebesgue measure in $\mathbb{R}^{n+1}$, as in [U] for analytic capacity.
@@
for some $f:E\to\mathbb{R}$ Borel function, and where $\mu:=\pazocal{H}^{n+1}_{p_{s}}|_{E_{p_{s}}}$ coincides, up to a constant (depending on $n$ and $s$), with the probability measure supported on $E_{p_{s}}$ that satisfies $\mu(Q_{j}^{k})=((\delta+1)\delta^{n})^{-k}$ for all $j,k$. By (4.8) $\mu$ (and also $|\nu|$) has upper $s$-parabolic growth of degree $n+1$. By analogous arguments to those of Lemma 4.4, there must exist a constant $\kappa(n,s)$ so that
@@
(4.11) $\geq f(\overline{x}_{0})-\varepsilon\frac{\mu(Q_{\overline{x}_{0}}^{k})}{\mu(Q_{j}^{h})}=f(\overline{x}_{0})-((\delta+1)\delta^{n})^{h-k}\varepsilon.$
@@
$\varepsilon<\frac{((\delta+1)\delta^{n})^{-m}}{2}f(\overline{x}_{0}),$
@@
Observe that $\text{dist}_{p_{s}}(\overline{z},Q_{\overline{z}}^{k}\setminus{Q_{\overline{z}}^{k+m}})\gtrsim\ell(Q_{\overline{z}}^{k+m})$ (with implicit constants depending on the non-self-intersection parameter $\delta=\delta(s)$), so we are able to estimate $|\pazocal{P}_{\nu^{-}}^{s}\chi_{Q_{\overline{z}}^{k}\setminus{Q_{\overline{z}}^{k+m}}}(\overline{z})|$ from above in the following way
@@
$\lesssim\frac{\nu^{-}(Q_{\overline{z}}^{k})}{\ell(Q_{\overline{z}}^{k+m})^{n+1}}\leq\varepsilon\frac{\mu(Q_{\overline{z}}^{k})}{\ell(Q_{\overline{z}}^{k+m})^{n+1}}=\varepsilon\frac{((\delta+1)\delta^{n})^{-k}}{((\delta+1)\delta^{n})^{-m}\ell(Q_{\overline{z}}^{k})^{n+1}}$
$=((\delta+1)\delta^{n})^{m}\varepsilon,$
@@
(4.15) $(\nabla_{x}P_{s})_{1}(\overline{z}-\bar{y})\geq 0\quad\mbox{ for all $\bar{y}\in Q^{k}_{\overline{z}}\setminus Q^{k+m}_{\overline{z}}$.}$
@@
$\widetilde{\pazocal{P}}^{s}_{\ast}\nu(\overline{z})\gtrsim(m-1)f(\overline{x}_{0})-C((\delta+1)\delta^{n})^{m}\varepsilon$
@@ 5. The non-comparability of $\Gamma_{\Theta}$ and $\gamma_{\Theta}^{1/2}$ in the plane
$|\langle T,\varphi\rangle|\lesssim\ell(Q)^{n+1},\qquad\text{for $Q\subset\mathbb{R}^{n+1}$ any $1$-parabolic cube and $\varphi$ admissible for $Q$}.$
@@
[…]
@@
We shall contradict the estimate presented in Lemma 5.3 by finding a point $\overline{x}_{0}\in E$ such that $\pazocal{T}^{\ast}\nu(\overline{x}_{0})$ is arbitrarily big. To this end, firstly, we properly choose such point by a similar argument to that of the proof of [MPrT, Theorem 6.3]. Pick $\overline{x}_{0}=(0,t_{0})\in E$ with $10<t_{0}<$ 1 a Lebesgue point for the density $f=\mathop{}\!\mathrm{d}\nu/\text{d}\pazocal{H}_{p_{1}}^{2}|_{E}$ satisfying $f(\overline{x}_{0})>0$, that can be done since $\nu(E)>0$. Hence, for $\varepsilon>0$ small enough, which will be fixed later on, there is an integer $k>0$ big enough so that if $B_{k,0}$ is the $1$-parabolic ball centered at $\overline{x}_{0}$ with radius $r(B_{k,0})=2^{-k}$,
@@
[…]

### v4

@@ 4.1. Existence of removable sets with positive $\pazocal{H}_{p_{s}}^{n+1}$ measure
Recall that for a given sequence $\lambda:=(\lambda_{k})_{k},\,0<\lambda_{k}<1/2$, we define its associated Cantor set $E\subset\mathbb{R}^{n+1}$ by the following algorithm: set $Q^{0}:=[0,1]^{n+1}$ the unit cube of $\mathbb{R}^{n+1}$ and consider $2^{n+1}$ disjoint (Euclidean) cubes inside $Q^{0}$ of side length $\ell_{1}:=\lambda_{1}$, with sides parallel to the coordinate axes and such that each cube contains a vertex of $Q^{0}$. Continue this same process now for each of the $2^{n+1}$ cubes from the previous step, but now using a contraction factor $\lambda_{2}$. That is, we end up with $2^{2(n+1)}$ cubes with side length $\ell_{2}:=\lambda_{1}\lambda_{2}$. Proceeding inductively we have that at the $k$-th step of the iteration we encounter $2^{k(n+1)}$ cubes, that we denote $Q_{j}^{k}$ for $1\leq j\leq 2^{k(n+1)}$, with side length $\ell_{k}:=\prod_{j=1}^{k}\lambda_{j}$. We will refer to them as cubes of the $k$-th generation. We define
@@
If $(E_{k})$ is the number of cubes of $E_{k}$, the above property followed, in essence, from
(4.5) $(E_{k})\,\ell_{k}^{n}=2^{k(n+1)}\,\ell_{k}^{n}=1,\;\;\forall k\geq 1.$
@@
Hence, it is clear that in order to obtain a potentially non-removable Cantor set $E$, one has to enlarge it. One way to do it (motivated by [MPrT, §6]) is as follows: let us fix $s\in(1/2,1]$ and choose what we call the non-self-intersection parameter $d\in\mathbb{Z}_{+}$, the minimum integer $d=d(s)\geq 2$ satisfying
$s>\frac{\log_{d}(d+1)}{2},\qquad\text{that is}\qquad d+1<d^{2s}.$
Let $Q^{0}:=[0,1]^{n+1}$ be the unit cube of $\mathbb{R}^{n+1}$ and consider $(d+1)d^{n}$ disjoint $s$-parabolic cubes $Q_{i}^{1},\,1\leq i\leq(d+1)d^{n}$, contained in $Q^{0}$, with sides parallel to the coordinate axes, side length $0<\lambda_{1}<1/d$, and disposed as follows: first, we consider the first $n$ intervals of the cartesian product $Q^{0}:=[0,1]^{n+1}$ (that is, those contained in spatial directions) and we divide, each one, into $d$ equal subintervals $I_{1},\ldots,I_{d}$. For each subinterval $I_{j}$, we contain another one $J_{j}$ of length $\lambda_{1}$. Now, we distribute $J_{1},\ldots,J_{d}$ in an equispaced way, fixing $J_{1}$ to start at 0 and $J_{d}$ to end at 1. More precisely, if for each interval $[0,1]$ we name
$l_{d}:=\frac{1-d\lambda_{1}}{d-1},\qquad J_{j}:=\big[(j-1)(\lambda_{1}+l_{d}),j\lambda_{1}+(j-1)l_{d}\big],\quad j=1,\ldots,d,$
we keep the following union of $d$ closed disjoint intervals of length $\lambda_{1}$
$T_{d}:=\bigcup_{j=1}^{d}J_{j}.$
Finally, for the remaining temporal interval $[0,1]$, we do the same splitting but in $d+1$ intervals of length $\lambda_{1}^{2s}$. That is, we name
$\widetilde{l}_{d}:=\frac{1-(d+1)\lambda_{1}^{2s}}{d},\qquad\widetilde{J}_{j}:=\big[(j-1)(\lambda_{1}^{2s}+\widetilde{l}_{d}),j\lambda_{1}^{2s}+(j-1)\widetilde{l}_{d}\big],\quad j=1,\ldots,d+1.$
Now we keep the union of $d+1$ closed disjoint intervals of length $\lambda_{1}^{2s}$
$\widetilde{T}_{d}:=\bigcup_{j=1}^{d+1}\widetilde{J}_{j}.$
From the above, we define the first generation of the Cantor set as $E_{1,p_{s}}:=(T_{d})^{n}\times\widetilde{T}_{d}$, that is conformed by $(d+1)d^{n}$ disjoint $s$-parabolic cubes (see Figure 1). This procedure continues inductively, i.e. the next generation $E_{2,p_{s}}$ will be the family of $(d+1)^{2}d^{2n}$ disjoint $s$-parabolic cubes of side length $\ell_{2}:=\lambda_{1}\lambda_{2}$, $\lambda_{2}<1/d$, obtained from applying the previous construction to each of the cubes of $E_{1,p_{s}}$. More generally, the $k$-th generation $E_{k,p_{s}}$ will be formed by $(d+1)^{k}d^{nk}$ disjoint $s$-parabolic cubes with side length $\ell_{k}:=\lambda_{1}\cdots\lambda_{k}$, $\lambda_{k}<1/d$, and with locations determined by the above iterative process. We name such cubes $Q^{k}_{j}$, with $j=1,\ldots,(d+1)^{k}d^{nk}$. The resulting $s$-parabolic Cantor set is
@@
(a) Depiction of $E_{1,p_{1}}$ with $d=2$
(b) Depiction of $E_{1,p_{2/3}}$ with $d=3$.
@@
This way, the critical value becomes $((d+1)d^{n})^{-1/(n+1)}<1/d$. For instance, if
(4.7) $\lambda_{j}:=((d+1)d^{n})^{-1/(n+1)},\qquad\text{for every }\,j,$
@@
$(E_{k,p_{s}})\cdot\ell_{k}^{n+1}=(d+1)^{k}d^{nk}\cdot\ell_{k}^{n+1}=1,\;\;\forall k\geq 1.$
Using the previous fact one can deduce $\pazocal{H}^{n+1}_{p_{s}}(E_{p_{s}})>0$. Indeed, consider the probability measure $\mu$ defined on $E_{p_{s}}$ such that for each generation $k$, $\mu(Q_{j}^{k}):=(d+1)^{-k}d^{-nk},\,1\leq j\leq(d+1)^{k}d^{nk}$. Let $Q$ be any $s$-parabolic cube, that we may assume to be contained in $Q^{0}$, and pick $k$ with the property $\ell_{k+1}\leq\ell(Q)\leq\ell_{k}$, so that $Q$ can meet, at most, $(d+1)d^{n}$ cubes $Q_{j}^{k}$. Thus $\mu(Q)\leq((d+1)d^{n})^{-(k-1)}$ and we deduce
(4.8) $\mu(Q)\lesssim\frac{1}{((d+1)d^{n})^{(k+1)}}=\ell_{k+1}^{n+1}\leq\ell(Q)^{n+1}$
@@
$\pazocal{H}^{n+1}_{p_{s},\varepsilon}(E)\leq\sum_{j=1}^{((d+1)d^{n})^{k}}\text{diam}_{p_{s}}(Q^{k}_{j})^{n+1}\simeq\ell_{k}^{n+1}\,((d+1)d^{n})^{k}=1.$
@@
Let us observe that as $s\to 1/2$, the value of $d$ grows arbitrarily. That is, as $s$ approaches the value $1/2$, the Cantor set $E_{p_{s}}$ becomes progressively more and more dense in the unit cube. This suggests that for $s=1/2$, that corresponds to a space-time usual Lipschitz condition in $\mathbb{R}^{n+1}$, the capacity one would obtain should be comparable to the Lebesgue measure in $\mathbb{R}^{n+1}$, as in [U] for analytic capacity.
@@
for some $f:E\to\mathbb{R}$ Borel function, and where $\mu:=\pazocal{H}^{n+1}_{p_{s}}|_{E_{p_{s}}}$ coincides, up to a constant (depending on $n$ and $s$), with the probability measure supported on $E_{p_{s}}$ that satisfies $\mu(Q_{j}^{k})=((d+1)d^{n})^{-k}$ for all $j,k$. By (4.8) $\mu$ (and also $|\nu|$) has upper $s$-parabolic growth of degree $n+1$. By analogous arguments to those of Lemma 4.4, there must exist a constant $\kappa(n,s)$ so that
@@
(4.11) $\geq f(\overline{x}_{0})-\varepsilon\frac{\mu(Q_{\overline{x}_{0}}^{k})}{\mu(Q_{j}^{h})}=f(\overline{x}_{0})-((d+1)d^{n})^{h-k}\varepsilon.$
@@
$\varepsilon<\frac{((d+1)d^{n})^{-m}}{2}f(\overline{x}_{0}),$
@@
Observe that $\text{dist}_{p_{s}}(\overline{z},Q_{\overline{z}}^{k}\setminus{Q_{\overline{z}}^{k+m}})\gtrsim\ell(Q_{\overline{z}}^{k+m})$ (with implicit constants depending on the non-self-intersection parameter $d=d(s)$), so we are able to estimate $|\pazocal{P}_{\nu^{-}}^{s}\chi_{Q_{\overline{z}}^{k}\setminus{Q_{\overline{z}}^{k+m}}}(\overline{z})|$ from above in the following way
@@
$\lesssim\frac{\nu^{-}(Q_{\overline{z}}^{k})}{\ell(Q_{\overline{z}}^{k+m})^{n+1}}\leq\varepsilon\frac{\mu(Q_{\overline{z}}^{k})}{\ell(Q_{\overline{z}}^{k+m})^{n+1}}=\varepsilon\frac{((d+1)d^{n})^{-k}}{((d+1)d^{n})^{-m}\ell(Q_{\overline{z}}^{k})^{n+1}}$
$=((d+1)d^{n})^{m}\varepsilon,$
@@
[…]
@@
$\widetilde{\pazocal{P}}^{s}_{\ast}\nu(\overline{z})\gtrsim(m-1)f(\overline{x}_{0})-C((d+1)d^{n})^{m}\varepsilon$
@@ 5. The non-comparability of $\Gamma_{\Theta}$ and $\gamma_{\Theta}^{1/2}$ in the plane
[…]
@@
$\pazocal{T}_{\mu}f(\overline{x}):=\int_{\mathbb{R}^{n+1}}(-\Delta)^{1/2}W(\overline{x}-\overline{y})f(\overline{y})\text{d}\mu(\overline{y}),\hskip 14.22636pt\overline{x}\notin\text{supp}(\mu),$
@@
We shall contradict the estimate presented in Lemma 5.3 by finding a point $\overline{x}_{0}\in E$ such that $\pazocal{T}^{\ast}\nu(\overline{x}_{0})$ is arbitrarily big. To this end, firstly, we properly choose such point by a similar argument to that of the proof of [MPrT, Theorem 6.3]. Pick $\overline{x}_{0}=(0,t_{0})\in E$ with $0<t_{0}<1$ a Lebesgue point for the density $f=\mathop{}\!\mathrm{d}\nu/\text{d}\pazocal{H}_{p_{1}}^{2}|_{E}$ satisfying $f(\overline{x}_{0})>0$, that can be done since $\nu(E)>0$. Hence, for $\varepsilon>0$ small enough, which will be fixed later on, there is an integer $k>0$ big enough so that if $B_{k,0}$ is the $1$-parabolic ball centered at $\overline{x}_{0}$ with radius $r(B_{k,0})=2^{-k}$,
@@
$|\pazocal{T}_{\nu^{-}}\chi_{B_{k,0}\setminus{B_{k+m,0}}}(\overline{x}_{0})|$ $\leq\int_{B_{k,0}\setminus{B_{k+m,0}}}|(-\Delta)^{1/2}W(\overline{x}_{0}-\overline{y})|d\nu^{-}(\overline{y})$